What volume in mL of 0.3000 M NaCl solution is required to produce 0.2250 moles of NaCl?


1
Expert's answer
2021-04-26T03:01:29-0400

Molar Concentration = number of moles/ volume


volume = number of moles/ molar concentration


volume = 0.2250/ 0.3000 = 0.75 L


volume required = 0.75 L = 0.75 L× 1000 mL/L = 750 mL